{ { { z/z+5 - 6/z-5 = 50/z^2-25 } } }
the LCD is (z+5)(z-5), but when I do it out I get a really jumbled answer that I know isn't correct.

Answer by venugopalramana(3286)

{ { { z/z+5 - 6/z-5 = 50/z^2-25 } } }
the LCD is (z+5)(z-5),
CORRECT...MULTIPLY WITH THAT THROUGH OUT
Z(Z-5)-6(Z+5)=50
Z^2-5Z-6Z-30=50
Z^2-11Z-80=0
Z^+5Z-16Z-80=0
Z(Z+5)-16(Z+5)=0
(Z-16)(Z+5)=0
Z-16=0...OR...Z=16.....OR
Z+5=....OR.....Z=-5
but when I do it out I get a really jumbled answer that I know isn't correct.